SUPPOSE THAT THE LENGTH OF A CERTAIN RECTANGLE IS THREE TIMES ITS WITDH. IF THE LENGTH IS DECREASED BY 4 INCHES AND THE WIDTH INCREASE BY 1 INCH, THE NEWLY FORMED RECTANGLE HAS AN AREA OF 40 SQUARE INCHES. FIND THE LENGTH AND WIDTH OF THE ORIGINAL RECTANGLE.
Answer by checkley77(12844)

L=3W-4
W=W+1
LW=40
(3w-4)(w+1)=40
3w^2-4w+3w-4=40
3w^2-w-4-40=0
3w^2-w-44=0
(3w+11)(w-4)=0
w-4=0
w=4 ans. for the width.
4+1=5 for the new width
L=3*4-4
L=12-4
L=8
PROOF:
5*8=40
40=40